# Sparse Feature Circuits: Discovering and Editing Interpretable Causal Graphs in Language Models

### v2

@@ Abstract
We introduce methods for discovering and applying sparse feature circuits. These are causally implicated subnetworks of human-interpretable features for explaining language model behaviors. Circuits identified in prior work consist of polysemantic and difficult-to-interpret units like attention heads or neurons, rendering them unsuitable for many downstream applications. In contrast, sparse feature circuits enable detailed understanding of unanticipated mechanisms. Because they are based on fine-grained units, sparse feature circuits are useful for downstream tasks: We introduce Shift, where we improve the generalization of a classifier by ablating features that a human judges to be task-irrelevant. Finally, we demonstrate an entirely unsupervised and scalable interpretability pipeline by discovering thousands of sparse feature circuits for automatically discovered model behaviors.
@@ 1 Introduction
The key challenge of interpretability research is to scalably explain the many unanticipated behaviors of neural networks (NNs). Much recent work explains NN behaviors in terms of coarse-grained model components, for example by implicating certain induction heads in in-context learning (Olsson et al., 2022) or MLP modules in factual recall (Meng et al., 2022; Geva et al., 2023; Nanda et al., 2023, inter alia). However, such components are generally polysemantic (Elhage et al., 2022) and hard to interpret, making it difficult to apply mechanistic insights to downstream applications. On the other hand, prior methods analyzing behaviors in terms of fine-grained units (Kim et al., 2018; Belinkov, 2022; Geiger et al., 2023; Zou et al., 2023) demand that researchers begin with the answer: they assume the existence of curated data that isolates the target behavior. Such approaches are not well-suited for discovering unanticipated mechanisms.
@@
We propose to explain model behaviors using fine-grained components that play narrow, interpretable roles. Doing so requires us to address two challenges: First, we must identify the correct fine-grained unit of analysis, since obvious choices like neurons [^1] We use “neuron” to refer to a basis-aligned direction in an LM’s latent space (not necessarily preceded by a nonlinearity). are rarely interpretable, and units discovered via supervised methods require pre-existing hypotheses. Second, we must address the scalability problem posed by searching for causal circuits over a large number of fine-grained units.
@@
We leverage recent progress in dictionary learning (Bricken et al., 2023; Cunningham et al., 2024) to tackle the first challenge. Namely, we train sparse autoencoders (SAEs) to identify directions in an LM’s latent space which represent human-interpretable features. Then, to solve the scalability challenge, we employ linear approximations (Sundararajan et al., 2017; Nanda, 2022; Syed et al., 2023) to efficiently identify SAE features which are most causally implicated in model behaviors, as well as connections between these features. The result is a sparse feature circuit which explains how model behaviors arise via interactions among fine-grained human-interpretable units.
@@
Sparse feature circuits can be productively used in downstream applications. We introduce a technique, Sparse Human-Interpretable Feature Trimming (Shift), which shifts the generalization of an LM classifier by surgically removing sensitivity to unintended signals—even without knowing what those signals are in advance. We demonstrate Shift by debiasing a classifier in a worst-case setting where an unintended signal (gender) is perfectly predictive of target labels (profession).
@@
Finally, we demonstrate our method’s scalability by automatically discovering thousands of LM behaviors with the clustering approach of Michaud et al. (2023), and then automatically discovering feature circuits for these behaviors.
@@
Figure 1: Overview. Given contrastive input pairs, classification data, or automatically discovered model behaviors, we discover circuits composed of human-interpretable sparse features to explain their underlying mechanisms. We then label each feature according to what it activates on or causes to happen. Finally, if desired, we can ablate spurious features out of the circuit to modify how the system generalizes.
@@
    A scalable method to discover sparse feature circuits. We validate our method by evaluating feature circuits on a suite of subject-verb agreement tasks.
@@
    Shift, a technique for removing sensitivity to unintended signals without disambiguating data.
@@
    A fully-unsupervised automatic feature circuit discovery pipeline that identifies circuits for thousands of automatically discovered LM behaviors.
@@ 2 Formulation
A fundamental challenge in NN interpretability is that individual neurons are rarely interpretable (Elhage et al., 2022). Recently, Cunningham et al. (2024); Bricken et al. (2023) have shown that sparse autoencoders (SAEs) can be used to identify interpretable directions. We closely follow Bricken et al. (2023) to train SAEs for attention outputs, [^2] In other words, the output of the attention layer’s out projection. MLP outputs, and residual stream activations for each layer of Pythia-70M (Biderman et al., 2023). Given an input activation ${\mathbf{x}}\in{\mathbb{R}}^{d_{\text{model}}}$ from one of these model components, the corresponding SAE computes a decomposition
@@
into an approximate reconstruction $\hat{{\mathbf{x}}}$ as a sparse sum of features ${\mathbf{v}}_{i}$ and an SAE error term ${\bm{\epsilon}}({\mathbf{x}})\in{\mathbb{R}}^{d_{\text{model}}}$. The features ${\mathbf{v}}_{i}\in{\mathbb{R}}^{d_{\text{model}}}$ are unit vectors, the feature activations $f_{i}({\mathbf{x}})\in{\mathbb{R}}$ are a sparse set of coefficients, ${\mathbf{b}}\in{\mathbb{R}}^{d_{\text{model}}}$ is a bias, and we take $d_{\text{SAE}}=64\cdot d_{\text{model}}$. The SAEs are trained to minimize an L2 reconstruction error and an L1 regularization term which promotes sparsity. Details about our SAEs and their training can be found in Appendix B.
[…]
@@
Let $m$ be a real-valued metric computed via a computation graph (e.g., a NN); let ${\mathbf{a}}\in{\mathbb{R}}^{d}$ represent a node in this graph. Following prior work (Vig et al., 2020; Finlayson et al., 2021), we quantify the importance of ${\mathbf{a}}$ on a pair of inputs $(x_{\text{clean}},x_{\text{patch}})$ via its indirect effect (IE; Pearl, 2001) on $m$:
@@
Here ${\mathbf{a}}_{\text{patch}}$ is the value that ${\mathbf{a}}$ takes in the computation of $m(x_{\text{patch}})$, and $m(x_{\text{clean}}|\text{do}({\mathbf{a}}={\mathbf{a}}_{\text{patch}}))$ denotes the value of $m$ when computing $m(x_{\text{clean}})$ but intervening in the computation of $m$ by manually setting ${\mathbf{a}}$ to ${\mathbf{a}}_{\text{patch}}$. For example, given inputs $x_{\text{clean}}=$“The teacher” and $x_{\text{patch}}=$“The teachers,” we have metric $m(x)=\log P(\text{``{\color[rgb]{0.88671875,0.12890625,0.2734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.88671875,0.12890625,0.2734375}are}''}|x)-\log P(\text{``{\color[rgb]{0.2421875,0.453125,0.8203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2421875,0.453125,0.8203125}is}''}|x)$ the log probability difference output by a LM. Then if ${\mathbf{a}}$ is the activation of a particular neuron, a large value of $\text{IE}(m;{\mathbf{a}};x_{\text{clean}},x_{\text{patch}})$ indicates that the neuron is highly influential on the model’s decision to output “is” vs. “are” on this pair of inputs.
We often want to compute IEs for a very large number of model components ${\mathbf{a}}$, which cannot be done efficiently with (2). We thus employ linear approximations to (2) that can be computed for many ${\mathbf{a}}$ in parallel. The simplest such approximation, attribution patching (Nanda, 2022; Syed et al., 2023; Kramár et al., 2024), employs a first-order Taylor expansion
@@
In practice, since we use small models, we can instead employ a more expensive but more accurate approximation based on integrated gradients (Sundararajan et al., 2017):
$\hat{\text{IE}}_{\text{ig}}(m;{\mathbf{a}};x_{\text{clean}},x_{\text{patch}})=\left(\sum_{\alpha}\left.\nabla_{{\mathbf{a}}}m\right|_{\alpha{\mathbf{a}}_{\text{clean}}+(1-\alpha){\mathbf{a}}_{\text{patch}}}\right)({\mathbf{a}}_{\text{patch}}-{\mathbf{a}}_{\text{clean}})$ (4)
where the sum in (4) ranges over $N=10$ equally-equally spaced $\alpha\in\{0,\tfrac{1}{N},\dots,\tfrac{N-1}{N}\}$. This cannot be done in parallel for two nodes when one is downstream of another, but can be done in parallel for arbitrarily many nodes which do not depend on each other. Thus the additional cost of computing $\hat{\text{IE}}_{\text{ig}}$ over $\hat{\text{IE}}_{\text{atp}}$ scales linearly in $N$ and the serial depth of $m$’s computation graph.
The above discussion applies to the setting where we have a pair of clean and patch inputs, and we would like to understand that effect of patching a particular node from its clean to patch values. But in some settings (§4, 5), we have only a single input $x$. In this case, we instead approximate the indirect effect $\text{IE}(m;{\mathbf{a}};x)=m(x|\text{do}({\mathbf{a}}={\mathbf{0}}))-m(x)$ of setting ${\mathbf{a}}$ to ${\mathbf{0}}$. We get the modified formulas for $\hat{\text{IE}}(m;{\mathbf{a}};x)$ from (3) and (4) by replacing ${\mathbf{a}}$ with ${\mathbf{0}}$.
@@ 3 Sparse Feature Circuit Discovery
In this section, we introduce sparse feature circuits, which are computational sub-graphs that explain model behaviors in terms of SAE features and error terms. We first explain our circuit discovery algorithm (§3.1). Then, we analyze circuits found with our technique to show that they are both more interpretable and more concise than circuits consisting of neurons (§3.2). Finally, we perform a case study of the circuits we discovered for two linguistic tasks related to subject-verb agreement (§3.3).
@@ 3.1 Method
Suppose we are given an LM $M$, SAEs for various submodules of $M$ (e.g., attention outputs, MLP outputs, and residual stream vectors, as in §2), a dataset $\mathcal{D}$ consisting either of contrastive pairs $(x_{\text{clean}},x_{\text{patch}})$ of inputs or of single inputs $x$, and a metric $m$ depending on $M$’s output when processing data from $\mathcal{D}$. For example, Figure 2 shows the case where $\mathcal{D}$ consists of pairs of inputs which differ in number, and $m$ is the log probability difference between $M$ outputting the verb form that is correct for the patch vs. clean input.
@@
Approximating the IE of each node. Let $\hat{\text{IE}}$ be one of $\hat{\text{IE}}_{\text{atp}}$ or $\hat{\text{IE}}_{\text{ig}}$ (see §2). Then for each node ${\mathbf{a}}$ in $G$ and input $x\sim\mathcal{D}$, we compute $\hat{\text{IE}}(m;{\mathbf{a}};x)$. We apply some choice of node threshold $T_{N}$ to select nodes with a large (absolute) IE.
Consistent with prior work (Nanda, 2022; Kramár et al., 2024), we find that $\hat{\text{IE}}_{\text{atp}}$ accurately estimates IEs for SAE features and SAE errors, with the exception of nodes in the layer 0 MLP and early residual stream layers, where $\hat{\text{IE}}_{\text{atp}}$ underestimates the true IE. We find that $\hat{\text{IE}}_{\text{ig}}$ significantly improves accuracy for these components, so we use it in our experiments below. See Appendix C for more information about linear approximation quality.
@@
Practical considerations. Various practical difficulties arise for efficiently computing the gradients needed by our method. We solve using a combination of stop gradients, pass-through gradients, and tricks for efficient Jacobian-vector product computation; see App. A.3.
3.2 Discovering and Evaluating Sparse Feature Circuits for Subject–Verb Agreement
@@
To evaluate our method, we discover sparse feature circuits (henceforth, feature circuits) for four variants of the subject-verb agreement task (Table 1). Specifically, we adapt data from Finlayson et al. (2021) to produce datasets consisting of contrastive pairs of inputs that differ only in the grammatical number of the subject; the model’s task is to choose the appropriate verb inflection. We use this data along with our SAEs from §2 to discover circuits for Pythia-70M.
@@
We evaluate circuits for interpretability, faithfulness, and completeness. For each criterion, we compare to neuron circuits discovered by applying our methods with neurons in place of sparse features; in this setting, there are no error terms ${\bm{\epsilon}}$.
@@
Interpretability. We asked human crowdworkers to rate the interpretability of random features, random neurons, features from our feature circuits, and neurons from our neuron circuits. Crowdworkers rate sparse features as significantly more interpretable than neurons, with features that participate in our circuits also being more interpretable than randomly sampled ones; see App. D.
@@
Figure 3: Faithfulness for circuits (a) and their complements (b), measured on held-out data. Faint lines correspond to the structures from Table 1, with the average in bold. The ideal faithfulness for circuits is 1, while the ideal score for their complements is 0.
@@
Faithfulness. Given a circuit $C$ and metric $m$, let $m(C)$ denote the average value of $m$ over $\mathcal{D}$ when running our model with all nodes outside of $C$ mean-ablated, i.e., set to their average value over data from $\mathcal{D}$ [^3] Following Wang et al. (2023), we ablate features by setting them to their mean position-specific values. We then measure faithfulness as $\frac{m(C)-m(\varnothing)}{m(M)-m(\varnothing)}$, where $\varnothing$ denotes the empty circuit and $M$ denotes the full model. Intuitively, this metric captures the proportion of the model’s performance our circuit explains, relative to mean ablating the full model (which represents the “prior” performance of the model when it is given information about the task, but not about specific inputs). We find that model components in early layers typically handle specific tokens; since the tokens appearing in our held-out evaluation set don’t necessarily align with those in our circuit discovery set, there isn’t prima facie reason for early circuit components to generalize. We thus measure the faithfulness of our circuits starting at layer 2.
@@
We plot faithfulness for feature circuits and neuron circuits after sweeping over node thresholds (Figure 3a). We find that small feature circuits explain a large proportion of model behavior: the majority of performance is explained by less than $100$ nodes. In contrast, around $1500$ neurons are required to explain half the performance. However, as SAE errors are high-dimensional and coarse-grained, they cannot be fairly compared to neurons; we thus also plot the faithfulness of feature circuits with all SAE error nodes removed, or with all attention and MLP error nodes removed. We find (unsurprisingly) that removing residual stream SAE errors severely disrupts the model and curtails its maximum performance. Removing MLP and attention errors is less disruptive.
@@
Completeness. Are there parts of the model behavior that our circuit fails to capture? We measure this as the faithfulness of the circuit’s complement $M\setminus C$ (Figure 3b). We observe that we can eliminate the model’s task performance by ablating only a few nodes from our feature circuits, and that this is true even when we leave all SAE errors in place. In contrast, it takes hundreds of neurons to acheive the same effect.
@@
3.3 Case study: Subject–verb agreement across a relative clause

### v3

@@ Abstract
We introduce methods for discovering and applying sparse feature circuits. These are causally implicated subnetworks of human-interpretable features for explaining language model behaviors. Circuits identified in prior work consist of polysemantic and difficult-to-interpret units like attention heads or neurons, rendering them unsuitable for many downstream applications. In contrast, sparse feature circuits enable detailed understanding of unanticipated mechanisms in neural networks. Because they are based on fine-grained units, sparse feature circuits are useful for downstream tasks: We introduce Shift, where we improve the generalization of a classifier by ablating features that a human judges to be task-irrelevant. Finally, we demonstrate an entirely unsupervised and scalable interpretability pipeline by discovering thousands of sparse feature circuits for automatically discovered model behaviors.
@@ 1 Introduction
The key challenge of interpretability research is to scalably explain the many unanticipated behaviors of neural networks (NNs). Much recent work explains NN behaviors in terms of coarse-grained model components, for example by implicating certain induction heads in in-context learning (Olsson et al., 2022) or MLP modules in factual recall (Meng et al., 2022; Geva et al., 2023; Nanda et al., 2023, inter alia). However, such components are generally polysemantic (Elhage et al., 2022) and hard to interpret, making it difficult to apply mechanistic insights to downstream applications. On the other hand, prior methods for analyzing behaviors in terms of fine-grained units (Kim et al., 2018; Belinkov, 2022; Geiger et al., 2023; Zou et al., 2023) attempt to fit model internals to researcher-specified mechanistic hypotheses using researcher-curated data. These approaches are not well-suited to the many cases where researchers cannot anticipate ahead of time how models internally implement their surprising behaviors.
@@
We propose to explain model behaviors using fine-grained components that play narrow, interpretable roles. Doing so requires us to address two challenges: First, we must identify an appropriate fine-grained unit of analysis, since obvious choices like neurons [^1] We use “neuron” to refer to a basis-aligned direction in an LM’s latent space (not necessarily preceded by a nonlinearity). are rarely interpretable, and units discovered via supervised methods like linear probing require pre-existing hypotheses (Mueller et al., 2024). Second, we must address the scalability problem posed by searching for causal circuits over a large number of fine-grained units.
@@
We leverage recent progress in dictionary learning for NN interpretability (Bricken et al., 2023; Cunningham et al., 2024) to tackle the first challenge. Namely, we use sparse autoencoders (SAEs) to identify directions in LM latent spaces which represent human-interpretable concepts. Then, to address the scalability challenge, we employ linear approximations (Sundararajan et al., 2017; Nanda, 2022; Syed et al., 2023) to efficiently identify SAE features which are most causally implicated in model behaviors, as well as connections between these features. The result is a sparse feature circuit which explains how model behaviors arise via interactions among fine-grained human-interpretable units.
@@
Sparse feature circuits can be productively used in downstream applications. We introduce a technique, Sparse Human-Interpretable Feature Trimming (Shift; §4), which shifts the generalization of an LM classifier by surgically removing sensitivity to unintended signals. Unlike previous work on spurious cue removal—which isolates spurious signals using disambiguating data—Shift identifies unintended signals using interpretability and human judgement. We thus showcase Shift by debiasing a classifier in a worst-case setting, where an unintended signal (gender) is perfectly predictive of target labels (profession).
@@
Finally, we demonstrate our method’s scalability by automatically discovering thousands of narrow LM behaviors—for example, predicting “to” as an infinitive object or predicting commas in dates—with the clustering approach of Michaud et al. (2023), and then automatically discovering feature circuits for these behaviors (§5).
@@
Figure 1: Overview. Given contrastive input pairs, classification data, or automatically discovered model behaviors, we discover circuits composed of human-interpretable sparse features to explain their underlying mechanisms. We then label each feature according to what it activates on or causes the model to predict. Finally, if desired, we can ablate spurious features out of the circuit to modify how the system generalizes.
@@
    A scalable method to discover sparse feature circuits. We validate our method by discovering and evaluating feature circuits on a suite of subject-verb agreement tasks.
@@
    Shift, a technique for removing a LM classifier’s sensitivity to unintended signals, even without data that isolate these signals.
@@
    A fully-unsupervised pipeline for computing feature circuits for thousands of automatically discovered LM behaviors, viewable at feature-circuits.xyz.
@@ 2 Formulation
A fundamental challenge in NN interpretability is that individual neurons are rarely interpretable (Elhage et al., 2022). Therefore, many interpretability researchers have recently turned to sparse autoencoders (SAEs), an unsupervised technique for identifying a large number of interpretable NN latents (Cunningham et al., 2024; Bricken et al., 2023; Templeton et al., 2024; Rajamanoharan et al., 2024a; b). Given a model component with latent space ${\mathbb{R}}^{d}$ and an activation ${\mathbf{x}}\in{\mathbb{R}}^{d}$, an SAE computes a decomposition
@@
into an approximate reconstruction $\hat{{\mathbf{x}}}$ as a sparse sum of features ${\mathbf{v}}_{i}$ and an SAE error term ${\bm{\epsilon}}({\mathbf{x}})\in{\mathbb{R}}^{d}$. Here $d_{\text{SAE}}$ is the width of the SAE, the features ${\mathbf{v}}_{i}\in{\mathbb{R}}^{d}$ are unit vectors, the feature activations $f_{i}({\mathbf{x}})\geq 0$ are a sparse set of coefficients, and ${\mathbf{b}}\in{\mathbb{R}}^{d}$ is a bias. SAEs are trained on an objective which promotes having a small reconstruction error $\|{\mathbf{x}}-\hat{{\mathbf{x}}}\|_{2}$ while using only a sparse set of feature activations $f_{i}({\mathbf{x}})$. Rather than discard the error terms ${\bm{\epsilon}}$ for the purposes of circuit discovery, our methods handle them gracefully by incorporating them into our sparse feature circuits; this gives a principled decomposition of model behaviors into contributions from interpretable features and error components not yet captured by our SAEs.
In this work, we leverage the following suites of SAEs:
  • A suite of SAEs we train for each sublayer (attention layer, MLP, residual stream, and embeddings) of Pythia-70M (Biderman et al., 2023). We closely follow Bricken et al. (2023), using a ReLU-linear encoder $f_{i}$ and sparse dimension $d_{\text{SAE}}=64\times d$ and training the SAEs to minimize a combination of an L2 reconstrution loss and L1 regularization term which promotes sparsity. Details about our Pythia SAEs and their training can be found in Appendix B.1.
  • The open source Gemma Scope SAEs (Lieberum et al., 2024) available for all sublayers (excluding embeddings) of the open-weights Gemma-2-2B model (Team et al., 2024). These SAEs use a Jump-ReLU-linear encoder and $d_{\text{SAE}}=8\times d$. Details about the Gemma Scope SAEs can be found in Appendix B.2.
Scalably training better SAEs is an active area of research, as illustrated by the ready availability of open-source SAEs (Gao et al., 2024; Lieberum et al., 2024; Lin & Bloom, 2023). Thus, our goal is to—given a suite of trained SAEs—scalably apply them to understand NN behaviors; we treat scaling the SAEs themselves as out-of-scope.
@@
Let $m$ be a real-valued metric computed via a computational graph (e.g., a NN); let ${\mathbf{a}}$ represent a node in this graph. Following prior work (Vig et al., 2020; Finlayson et al., 2021), we quantify the importance of ${\mathbf{a}}$ on a pair of inputs $(x_{\text{clean}},x_{\text{patch}})$ via its indirect effect (IE; Pearl, 2001) on $m$:
@@
Here, ${\mathbf{a}}_{\text{patch}}$ is the value that ${\mathbf{a}}$ takes in the computation of $m(x_{\text{patch}})$, and $m(x_{\text{clean}}|\text{do}({\mathbf{a}}={\mathbf{a}}_{\text{patch}}))$ denotes the value of $m$ when computing $m(x_{\text{clean}})$ but intervening in the computation of $m$ by manually setting ${\mathbf{a}}$ to ${\mathbf{a}}_{\text{patch}}$. For example, given inputs $x_{\text{clean}}=$“The teacher” and $x_{\text{patch}}=$“The teachers,” we have metric $m(x)=\log P(\text{``{\color[rgb]{0.88671875,0.12890625,0.2734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.88671875,0.12890625,0.2734375}are}''}|x)-\log P(\text{``{\color[rgb]{0.2421875,0.453125,0.8203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2421875,0.453125,0.8203125}is}''}|x)$, the log probability difference output by the LM. Then if ${\mathbf{a}}$ is the activation of a particular neuron, a large value of $\text{IE}(m;{\mathbf{a}};x_{\text{clean}},x_{\text{patch}})$ indicates that the neuron is highly influential on the model’s decision to output “is” vs. “are” on this pair of inputs.
We often want to compute IEs for a very large number of model components ${\mathbf{a}}\in{\mathbb{R}}^{d}$, which cannot be done efficiently with (2). We thus employ linear approximations to (2) that can be computed for many ${\mathbf{a}}$ in parallel. The simplest such approximation, attribution patching (Nanda, 2022; Syed et al., 2023; Kramár et al., 2024), employs a first-order Taylor expansion
@@
To improve the quality of the approximation, we can instead employ a more expensive but more accurate approximation based on integrated gradients (Sundararajan et al., 2017; Hanna et al., 2024):
$\hat{\text{IE}}_{\text{ig}}(m;{\mathbf{a}};x_{\text{clean}},x_{\text{patch}})=\frac{1}{N}\left(\sum_{\alpha}\left.\nabla_{{\mathbf{a}}}m\right|_{\alpha{\mathbf{a}}_{\text{clean}}+(1-\alpha){\mathbf{a}}_{\text{patch}}}\right)({\mathbf{a}}_{\text{patch}}-{\mathbf{a}}_{\text{clean}})$ (4)
where the sum in (4) ranges over $N=10$ equally-spaced $\alpha\in\{0,\tfrac{1}{N},\dots,\tfrac{N-1}{N}\}$. This cannot be done in parallel for two nodes when one is downstream of another, but can be done in parallel for arbitrarily many nodes which do not depend on each other. Thus the additional cost of computing $\hat{\text{IE}}_{\text{ig}}$ over $\hat{\text{IE}}_{\text{atp}}$ scales linearly in $N$ and the serial depth of $m$’s computation graph.
The above discussion applies to the setting where we have a pair of clean and patch inputs, and we would like to understand that effect of patching a particular node from its clean to patch values. But in some settings (e.g., §4, 5), we have only a single input $x$. In this case, we instead use a zero-ablation, using the indirect effect $\text{IE}(m;{\mathbf{a}};x)=m(x|\text{do}({\mathbf{a}}={\mathbf{0}}))-m(x)$ from setting ${\mathbf{a}}$ to ${\mathbf{0}}$. We get the modified formulas for $\hat{\text{IE}}(m;{\mathbf{a}};x)$ from (3) and (4) by replacing ${\mathbf{a}}$ with ${\mathbf{0}}$.
@@ 3.1 Method
Suppose we are given an LM $M$, SAEs for various submodules of $M$ (e.g., attention outputs, MLP outputs, and residual stream vectors, as in §2), a dataset $\mathcal{D}$ consisting either of contrastive pairs $(x_{\text{clean}},x_{\text{patch}})$ of inputs or of single inputs $x$, and a metric $m$ that depends on $M$’s output when processing data from $\mathcal{D}$. For example, Figure 2 shows the case where $\mathcal{D}$ consists of pairs of inputs which differ in number, and $m$ is the log probability difference between $M$ outputting the verb form that is correct for the patch vs. clean input.
@@
Approximating the IE of each node. Let $\hat{\text{IE}}$ be one of $\hat{\text{IE}}_{\text{atp}}$ or $\hat{\text{IE}}_{\text{ig}}$ (see §2). Then for each node ${\mathbf{a}}$ in $G$ and input $x\sim\mathcal{D}$, we compute $\hat{\text{IE}}(m;{\mathbf{a}};x)$; we then average over $x\sim\mathcal{D}$ to produce a score $\hat{\text{IE}}(m;{\mathbf{a}})$ and filter for nodes with $|\hat{\text{IE}}(m;{\mathbf{a}})|>T_{N}$ for some choice $T_{N}$ of node threshold.
Consistent with prior work (Nanda, 2022; Kramár et al., 2024), we find that $\hat{\text{IE}}_{\text{atp}}$ accurately estimates IEs for SAE features and SAE errors, with the exception of nodes in the layer 0 MLP and early residual stream layers, where $\hat{\text{IE}}_{\text{atp}}$ underestimates the true IE. We find that $\hat{\text{IE}}_{\text{ig}}$ significantly improves accuracy over $\hat{\text{IE}}_{\text{atp}}$ for these components, so we use it in our experiments below. See Appendix H for more information about linear approximation quality.
@@
Practical considerations. Various practical difficulties arise for efficiently computing the gradients needed by our method. We solve these using a combination of stop gradients, pass-through gradients, and tricks for efficient Jacobian-vector product computation; see App. A.3.
3.2 Discovering and Evaluating Sparse Feature Circuits for Subject-Verb Agreement
@@
To evaluate our method, we discover sparse feature circuits (henceforth, feature circuits) on Pythia-70M and Gemma-2-2B for four variants of the subject-verb agreement task (Table 1). Specifically, we adapt data from Finlayson et al. (2021) to produce datasets consisting of contrastive pairs of inputs that differ only in the grammatical number of the subject; the model’s task is to choose the appropriate verb inflection.
@@
We evaluate circuits for interpretability, faithfulness, and completeness. For each criterion, we compare to neuron circuits discovered by applying our methods with neurons in place of sparse features; in this setting, there are no error terms ${\bm{\epsilon}}$. When evaluating feature circuits for faithfuless and completeness, we use a test split of our dataset, consisting of contrastive pairs not used to discover the circuit.
@@
Interpretability. For Pythia SAEs, we asked human crowdworkers to rate the interpretability of random features, random neurons, features from our feature circuits, and neurons from our neuron circuits. Crowdworkers rated sparse features as significantly more interpretable than neurons, with features that participate in our circuits also being more interpretable than randomly sampled ones (App. F). This replicates prior findings that SAE features are substantially more interpretable than neurons (Bricken et al., 2023). For Gemma-2 SAEs, we refer the reader to Lieberum et al. (2024), which finds the interpretability of these SAEs’ features to be on par with those trained via other state-of-the-art techniques.
@@
Figure 3: Faithfulness and completeness scores for circuits, measured on held-out data. Faint lines correspond to the structures from Table 1, with the average across structures in bold. The ideal faithfulness for circuits is 1, while the ideal completeness is 0.
@@
Faithfulness. Given a circuit $C$ and metric $m$, let $m(C)$ denote the average value of $m$ over inputs from $\mathcal{D}$ when running our model with all nodes outside of $C$ mean-ablated, i.e., set to their average value over data from $\mathcal{D}$ [^2] Following Wang et al. (2023), we ablate features by setting them to their mean position-specific values. We then measure the faithfulness of a circuit as $\frac{m(C)-m(\varnothing)}{m(M)-m(\varnothing)}$, where $\varnothing$ denotes the empty circuit and $M$ denotes the full model. Intuitively, this metric captures the proportion of the model’s performance our circuit explains, relative to mean ablating the full model (which represents the “prior” performance of the model when it is given information about the task, but not about specific inputs).
We find that components in early model layers are typically involved in processing specific tokens. In practice, the inputs in the train split of our dataset (used to discover the circuit) and the test split (for evaluation) do not contain identical tokens, making it difficult to evaluate the quality of early segments of our circuit. Thus, we ignore the first $\nicefrac{{1}}{{3}}$ of our circuit, and only evaluate the latter $\nicefrac{{2}}{{3}}$.
@@
We plot faithfulness for feature circuits and neuron circuits after sweeping over node thresholds $T_{N}$ (Fig. 3). We find that small feature circuits explain a large proportion of model behavior: the majority of performance in Pythia-70M, resp. Gemma-2-2B is explained by only $100$, resp. $500$ nodes. In contrast, around $1500$, resp. $50000$ neurons are required to explain half the performance. However, as SAE error nodes are high-dimensional and coarse-grained, they cannot be fairly compared to neurons; we thus also plot the faithfulness of feature circuits with all SAE error nodes removed, or with all attention and MLP error nodes removed. Unsurprisingly, we find that removing residual stream SAE error nodes severely disrupts the model and curtails its maximum performance; removing MLP and attention error nodes is less disruptive.
@@
Completeness. Are there parts of the model behavior that our circuit fails to capture? We measure this as the faithfulness of the circuit’s complement $M\setminus C$ (Fig. 3). We observe that we can eliminate the model’s task performance by ablating only a few nodes from our feature circuits, and that this is true even when we leave all SAE errors in place. In contrast, it takes hundreds (for Pythia) or thousands (for Gemma) of neurons to achieve the same effect.
@@
3.3 Case study: Subject-verb agreement across a relative clause